Can you go step by step on this problem Im having trouble formulating the necessary formulas.
The Petersons drove to a friends house at 80km/h. On the return trip, rain reduced their speed by 10km/h. If their total driving time was 3h, find the total distance driven.
Thanks

Answer by kapilsinghi(68)

let the distance one way = x
time taken to go to friends place = x/80
time taken on way back = x/70
total time = x/80 + x/70 = 3
(70x + 80x)/(70*80) = 3
150x = 3*70*80
x = 3*70*80/150
x = 112 Km
total distance = 224 Km
kapilsinghi123@gmail.com